# Unveiling the electronic structure of the charge density wave and topological semimetal TaTe4 through high-field magnetotransport measurements

### v1

@@ Acknowledgements.
Data availability:
@@
Appendix A Effective masses
@@
Using this approach, and following the magnetic field orientations defined for configuration A in previous sections, we obtained $m^{cyc}_{\beta-0\textdegree}=(0.30\pm 0.03)m_{e}$ for the lower frequency branch of pocket $\beta$ at 0 and $m^{cyc}_{\beta-45\textdegree}=(0.32\pm 0.03)m_{e}$ at 45 in configuration A (figs. 6(a,b)), which suggests that the mass of charge carriers in these pockets is highly isotropic along this rotation plane.
Figure 6: (Color online) Temperature dependence of the amplitude of the oscillating component of magnetoresistance for TaTe4 for some dominant pockets and magnetic field orientations in Configuration A. Columns (a) and (b) show the oscillatory magnetoresistance data and corresponding LK fits for pocket $\beta$ at 0° and 45° respectively, while columns (c) and (d) are structured similarly but for pocket $\alpha$. The top row displays the oscillatory data filtered for each pocket at different magnetic fields and temperatures, the middle row shows the LK fit and the cyclotron mass associated to that fit of the data at the top row, and the bottom row shows the LK fit for a single magnetic field intensity and different temperatures. The middle column plots show the fits of the data to the LK formula of the selected pocket. The values of cyclotron effective masses are obtained from this fit. The right column shows the LK for a fixed magnetic field value, at different temperatures.
@@
On the other hand, pocket $\alpha$ seems to be anisotropic in terms of effective masses and shape (figs. 6(c,d)). The effective mass of the charge carriers smoothly changes with angle for pocket $\alpha$, with $m^{cyc}_{\alpha-45\textdegree}=(0.21\pm 0.03)m_{e}$ at 0 and $m^{cycl}_{\alpha-0\textdegree}=(0.10\pm 0.02)m_{e}$ at 45 in configuration A, indicating a high mass anisotropy for this pocket. Additionally, figure 6(c,ii) shows that the oscillation frequency of pocket $\alpha$ at 0 is $f_{min}=$44 T, whereas at 45 this value is $f_{45}=$32 T. Then, $K=(f_{45}/f_{min})^{2}=1.89$, which is similar to the effective mass anisotropy, $m^{cyc}_{\alpha-45\textdegree}/m^{cyc}_{\alpha-0\textdegree}=2.10\pm 0.72$.

### v2

@@ Acknowledgements.
Author Contributions
D.S-V., J.R-C and P.G-G. did the single crystal sinthesis. D.S-V., J.R-C, P.G-G., E.H-V., J.A.G. and S.C took the high field transport data. B.U., R.G-H and A.C.G-C. did the DFT calculations. All authors participated in the data analysis, discussion and manuscript writing and revision.
Data availability
@@
Supplementary Information
S.I Effective masses
@@
Using this approach, and following the magnetic field orientations defined for configuration A in previous sections, we obtained $m^{cyc}_{\beta-0\textdegree}=(0.30\pm 0.03)m_{e}$ for the lower frequency branch of pocket $\beta$ at 0 and $m^{cyc}_{\beta-45\textdegree}=(0.32\pm 0.03)m_{e}$ at 45 in configuration A (figs. S1(a,b)), which suggests that the mass of charge carriers in these pockets is highly isotropic along this rotation plane.
Figure S1: (Color online) Temperature dependence of the amplitude of the oscillating component of magnetoresistance for TaTe4 for some dominant pockets and magnetic field orientations in Configuration A. Columns (a) and (b) show the oscillatory magnetoresistance data and corresponding LK fits for pocket $\beta$ at 0° and 45° respectively, while columns (c) and (d) are structured similarly but for pocket $\alpha$. The top row displays the oscillatory data filtered for each pocket at different magnetic fields and temperatures, the middle row shows the LK fit and the cyclotron mass associated to that fit of the data at the top row, and the bottom row shows the LK fit for a single magnetic field intensity and different temperatures. The middle column plots show the fits of the data to the LK formula of the selected pocket. The values of cyclotron effective masses are obtained from this fit. The right column shows the LK for a fixed magnetic field value, at different temperatures.
@@
On the other hand, pocket $\alpha$ seems to be anisotropic in terms of effective masses and shape (figs. S1(c,d)). The effective mass of the charge carriers smoothly changes with angle for pocket $\alpha$, with $m^{cyc}_{\alpha-45\textdegree}=(0.21\pm 0.03)m_{e}$ at 0 and $m^{cycl}_{\alpha-0\textdegree}=(0.10\pm 0.02)m_{e}$ at 45 in configuration A, indicating a high mass anisotropy for this pocket. Additionally, figure S1(c,ii) shows that the oscillation frequency of pocket $\alpha$ at 0 is $f_{min}=$44 T, whereas at 45 this value is $f_{45}=$32 T. Then, $K=(f_{45}/f_{min})^{2}=1.89$, which is similar to the effective mass anisotropy, $m^{cyc}_{\alpha-45\textdegree}/m^{cyc}_{\alpha-0\textdegree}=2.10\pm 0.72$.